OpenStudy (anonymous):

the spinner on a wheel of fortune can land with an equal chance on any one of ten regions. three regions are red, four are blue, two are yellow, and one is green. a player wins $4 if the spinner stops on red and $2 if it stops on green. the player loses $2 if it stop on blue and $3 if it stops on yellow. what is the expected value? what does this value mean?

OpenStudy (kropot72):

Do you know the probability of the spinner landing on one of the 3 red regions?

OpenStudy (anonymous):

wouldn't it be 3 out of 10 ?

OpenStudy (kropot72):

Nearly right. P(R) = 3/10 = 0.3. Can you find P(B), P(Y) and P(G)?

OpenStudy (anonymous):

blue would be .4 yellow .2 and green .1 ?

OpenStudy (kropot72):

Correct! Now we need to set up a probability distribution table: Let the dollar profit be X.  Can you follow so far?

OpenStudy (kropot72):

@swat05 Are you there?

OpenStudy (kropot72):

E(X) = (4 * 0.3) + (-2 * 0.4) + (-3 * 0.2) + (2 * 0.1) = 0